If the market demand decreases by 50% what will be the new clearing price and


quantity for this market?

1
Expert's answer
2022-02-20T16:01:26-0500

"NCP=(1-\frac{50\%}{100\%})P=(1-0.5)P=0.5P,"

"Q=\frac 1{0.5}=2."